A spring has a spring constant of 450 N/m. How much must this spring be stretched to

store 49 J of potential energy?


1
Expert's answer
2021-05-27T17:48:21-0400

The potential energy of the spring is given as follows:


"E = \dfrac{kx^2}{2}"

where "k = 450N\/m" is the spring constant, and "x" the spring elongation. Expressing "x" and substituting "E = 49J", obtain:


"x = \sqrt{\dfrac{2E}{k}} = \sqrt{\dfrac{2\cdot 49}{450}} \approx 0.47m"

Answer. 0.47 m.